OpenStudy (vera_ewing):

What is the percentage yield of ammonia if 358 g N2 react with excess hydrogen to produce 357 g NH3? 3H2 + N2 2NH3

OpenStudy (unklerhaukus):

\[3\text H_2 + \text N_2 = 2\text{NH}_3\]

OpenStudy (unklerhaukus):

How many moles is 358 g of N_2 ?

OpenStudy (vera_ewing):

A. 82.1% B. 60.8% C. 50.0% D. 99.7%

OpenStudy (vera_ewing):

I know it's not D

OpenStudy (unklerhaukus):

the number of moles \(n\), is equal to the mass \(m\) divided by the molecular mass \(M\) \[n = m/M\]

OpenStudy (unklerhaukus):

do you know the molecular mass of N_2?

OpenStudy (vera_ewing):

no

OpenStudy (unklerhaukus):

a periodic table table will tell you the atomic mass of N, multiply this by two to get the molecular mass of N_2

OpenStudy (vera_ewing):

so the molecular mass of N is 14.01 right?

OpenStudy (vera_ewing):

so when multiplied by two, i get 28.02

OpenStudy (unklerhaukus):

good, so how many moles is 358 g of N_2 ?

OpenStudy (vera_ewing):

how do i calculate that?

OpenStudy (vera_ewing):

oh the answer is A right?

OpenStudy (unklerhaukus):

\[m= 358 [\text g]\\~\\M= 28.02[\text{g/mol}]\] \[n = m/M\]

OpenStudy (unklerhaukus):

i cannot predict the correct option without the working out

OpenStudy (unklerhaukus):

How many moles of N_2 are there?

OpenStudy (vera_ewing):

n=12.77

OpenStudy (unklerhaukus):

good

OpenStudy (vera_ewing):

now what?

OpenStudy (unklerhaukus):

Now, how many moles is 357 [g] of NH_3?

OpenStudy (unklerhaukus):

use \[M_{\text{NH}_3} = M_\text N+3M_\text H\]

OpenStudy (unklerhaukus):

\[M_\text N = 14.01\\~\\ M_\text H = 1.01\]

OpenStudy (vera_ewing):

MNH3=17.04

OpenStudy (unklerhaukus):

right, so how many moles?

OpenStudy (vera_ewing):

17.04? or is there more to do?

OpenStudy (unklerhaukus):

M = 17.04 [g/mol] is the molar mass there is a mass m = 357 g how many moles? n = m/M

OpenStudy (vera_ewing):

n=20.9

OpenStudy (unklerhaukus):

careful with rounding n = 357/17.04 = 20.9507 = 21.0 (3 sig.fig's)

OpenStudy (vera_ewing):

oops. i knew that :') now what?

OpenStudy (unklerhaukus):

Let's look at the equation again \[3\text H_2 + \text N_2 = 2\text{NH}_3\] We found (above) that 12.77 moles of N_2 were used How many moles of \(\text{NH}_3\) would we expect if the reaction reached completion?

OpenStudy (vera_ewing):

60 mol?

OpenStudy (vera_ewing):

@UnkleRhaukus

OpenStudy (unklerhaukus):

According to the equation, each mole of N_2 should form 2 moles of NH_3, (at completion)

OpenStudy (unklerhaukus):

so 12.77 moles of N_2 should form ? moles of NH_3

OpenStudy (vera_ewing):

24.154 moles?

OpenStudy (unklerhaukus):

try again 12.77 x 2 =

OpenStudy (vera_ewing):

25.54

OpenStudy (unklerhaukus):

yeah, lolz

OpenStudy (unklerhaukus):

OK, so we expected 25.54 moles of NH_3 product but only got 21.0 moles

OpenStudy (vera_ewing):

so the answer is A

OpenStudy (vera_ewing):

82.1%

OpenStudy (unklerhaukus):

the percentage yield is \[Y =\frac{\text{yield}}{\text{expected yield}}\]

OpenStudy (vera_ewing):

right?

OpenStudy (unklerhaukus):

yeah about \(82\%\)

OpenStudy (unklerhaukus):

Does it all make sense now?

OpenStudy (vera_ewing):

@UnkleRhaukus yeah it does. sorry for taking so long to respond. i had to eat dinner.